• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kai Pan

    Log entry time 14:56:16 on October 02, 2009

    Entry number 292747

    keyword=WAC report

    All of the following info has already been told to Dustin.

    1. #12861~#12869 are still pedstal runs and test runs lasting 1 or 2 minutes.

    2. #12870~#12875 are "Helicity Error" runs. I think we can still get those data back through software modification. However, I think it's better to let Happex expert to do this job. I still added these runs into slug as "production=0".

    3. #12876~#12880 are good production runs, however the Lumi RMS are all larger than 1200, and close 1300. Detector asymmetry width are normal. (#12878 was just running for 2 minutes, as the shift worker stopped it to check the spot++, set this run as "production=0")

    4. #12881~#12883 beam down, added as "production=0"

    5. #12884~#12885 beam back, and are good runs. All Lumi RMS recover normal which is less than 1000.

    6. Just find that the synchronization histogram is weird, looks like injector part has problem.

    A copy of this log entry has been emailed to: paschke,mercado,mcnulty,pkpan,mfriend